OpenStudy (idealist10):

How to integrate xe^(5/4*x^4) dx?

OpenStudy (idealist10):

\[xe ^{\frac{ 5 }{ 4 }x^4}dx\]

OpenStudy (faiqraees):

I guess integration by parts

OpenStudy (idealist10):

@Kainui

OpenStudy (kainui):

\[\frac{\sqrt{5}}{2}x^2=u\]\[\sqrt{5}x dx = du\] \[\int x e^{\frac{5}{4}x^4}dx = \frac{1}{\sqrt{5}} \int e^{u^2}du\] Not looking too good unless we have some bounds

OpenStudy (kainui):

Power series solution is super easy though! \[xe^{\frac{5}{4}x^4} = x \sum_{n=0}^\infty \frac{\left(\frac{5}{4}x^4\right)^n}{n!} = \sum_{n=0}^\infty \left(\frac{5}{4}\right)^n\frac{x^{4n+1}}{n!}\] Now it's just a polynomial, which we can integrate: \[\int \sum_{n=0}^\infty \left(\frac{5}{4}\right)^n\frac{x^{4n+1}}{n!} dx = \sum_{n=0}^\infty \int \left(\frac{5}{4}\right)^n\frac{x^{4n+1}}{n!}dx\]\[ = C+ \sum_{n=0}^\infty \left(\frac{5}{4}\right)^n \frac{1}{4n+2}\frac{x^{4n+2}}{n!}\]